US Steel is trading at $16.63 per share. The 16-strike call for June 2019 is quoted at $1.47 ask. What percent of the total option price is intrinsic value?
Answer #1

Intrinsic value of call=MAX(16.63-16,0)=0.63

% of total option value as intrinsic value=0.63/1.47=42.86%